b. Suppose a manufacturer of tractors secures a sale to a Chinese company of 240 million USD for delivery in 45 days. If market interest in China is 6.125% and market interest in the US is 3.2%, spot rate is RMB 6.831=1USD, calculate the expected forward rate and rate of appreciation/depreciation at the time of delivery. Show how the manufacturer can use a forward market hedge to lock in his/her profit.
Answer #1

Particulars

USD

Chinese RMB

Interest Rate

3.2%

6.125%

For 45 Days Interest Rates(365 Days taken)(3.2*45/365:6.125*45/365)

0.3945%

0.7551%

Ans:

Sport Rate 1 USD= 6.831 RMB

Forward Rate=1.003945=6.88258

             * 1.003945=(1+0.3945%)

             ** 6.88258=(6.831+0.7551%)

Therefore 1 dollar=6.8555

(6.88258/1.003945*1)

Expected rate of appreciation for USD at the time of Delivery=Forward rate-Spot rate/Spot rate*100

=6.8555-6.831/6.831*100=0.3586%

Therefore he is going to receive the dollars and dollars is going to be appreciated i.e he will get more units of their home currency due to home currency is going to week and dollar is going to up by 0.3586% , Therefore he need to take any forward hedging strategy

Forward hedging strategy is required when he feels that the home is going to be strong which results in he will get less number units of home currency when home currency is strong ,he can short sell dollars by putting 45 days time.